### v1

# Molecular Identification and Peak Assignment: Leveraging Multi-Level Multimodal Alignment on NMR

Abstract
Nuclear magnetic resonance (NMR) spectroscopy plays an essential role across various scientific disciplines, providing valuable insights into molecular dynamics and interactions. Despite the promise of AI-enhanced NMR prediction models, challenges persist in the interpretation of spectra for tasks such as molecular retrieval, isomer recognition, and peak assignment. In response, this paper introduces Multi-Level Multimodal Alignment with Knowledge-Guided Instance-Wise Discrimination (K-M3AID) to establish meaningful correspondences between two heterogeneous modalities: molecular graphs (structures) and NMR spectra. In particular, K-M3AID employs a dual-coordinated contrastive learning architecture, and incorporates a graph-level alignment module, a node-level alignment module, and a communication channel. Notably, the framework introduces knowledge-guided instance-wise discrimination into contrastive learning within the node-level alignment module, significantly enhancing accuracy in cross-modal alignment. Additionally, K-M3AID showcases its capability of meta-learning by demonstrating that skills acquired during node-level alignment positively impact graph-level alignment. Empirical validation underscores K-M3AID’s effectiveness in addressing multiple zero-shot tasks, offering a promising solution to bridge the gap between structural information and spectral data in complex NMR scenarios.
1 Introduction
Nuclear magnetic resonance (NMR) spectroscopy has found a board applications in various scientific domains, such as organic chemistry, pharmaceuticals, biochemistry, food science, and material science, by providing insights into molecular dynamics and interactions (Gunther & Gunther (1994); Claridge (2016); Yu et al. (2021)). The positions and shapes of signal peaks in NMR spectra, influenced by through-bond and through-space interactions, serve as ”fingerprints” for deducing atomic connectivity, relative stereochemistry, and conformations. The conventional approach for elucidating molecular structures and attributing peaks has long relied on manual determination by organic chemists (Guan et al. (2021)). However, the interpretation of NMR spectra is not straightforward, particularly when dealing with isomers and complex molecules consisting of multiple stereogentic (chiral) centers (Wu et al. (2023); Chhetri et al. (2018)). Even an expert chemist may encounter significant difficulties in accurately assigning isomeric compounds with extremely similar NMR spectra due to this complexity (Nicolaou & Snyder (2005)).
Recent advancements in AI-enhanced NMR prediction models show promise in generating spectra from candidate structures (Chen et al. (2020); Jonas et al. (2022); Kuhn (2022)). However, real-world scenarios often demand spectral interpretation before detailed structural information is available. Ideally, the interpretation of spectra should transition from spectral data to structural elucidation. Consequently, these prediction models are not applicable in highly demanding real-world tasks, such as molecular retrieval, candidate ranking, and peak assignment (see Figure 1.a). In scenarios with limited knowledge of an unknown compound, such as identifying a compound from a plant, molecular retrieval is crucial, as the retrieved analogs can aid in structural hypothesis formulation. In cases where pre-existing knowledge is available, such as identifying reaction products, the quantitative ranking of candidate possibilities becomes essential. Peak assignment represents the decisive stage in the structural identification process, providing detailed insights into the local atomic context. These prediction models fall short in identifying compounds without prior knowledge, quantitatively assessing candidates (especially in isomer recognition), and guiding accurate peak assignments. How might one employ or interrogate these advanced AI methodologies to enhance molecular retrieval, candidate ranking, and peak assignment in NMR spectra interpretation?
NMR spectroscopy can be viewed as a modality, capturing molecular structural and electronic details through an NMR spectrometer and expressing this information in sequence format (chemical shifts). On the other hand, these structural and electronic information can also be presented in a molecular graph, which encapsulates molecular attributes in a graph-based format. The tasks of determining molecular structure and assigning peaks require clear correspondence across these two heterogeneous modalities: NMR spectra and molecular graphs. This imperative is met by multimodal learning, particularly multimodal alignment (see fundamentals of multimodal alignment in Preliminary  2). However, multimodal alignment commonly encounters challenges in addressing data heterogeneity, semantic gaps among modalities, collecting paired data, designing effective cross-modal attention mechanisms, and formulating appropriate evaluation metrics (Liang et al. (2023); Jabeen et al. (2023)).
Molecules is composed of atoms, and each molecule, with its unique arrangement of atoms and bonding patterns, gives rise to a distinct spectrum. Consequently, it is impractical to incorporate all molecules and their spectra in the training dataset. One of the potential solutions is to take into account the constrained possibilities of atomic diversity for interpreting spectra. This, in turn, motivates us to leverage the power of AI for interpreting spectra at the atomic level, establishing a solid foundation for molecular identification. This approach aligns tightly with the core principles of meta-learning (see principles of meta-learning in Preliminary  2). However, meta-learning typically concerns in meta-overfitting, distance evaluation metrics, model interpretablity, transferability and generalization (Huisman et al. (2021); Feng et al. (2022); Ma et al. (2022); Vettoruzzo et al. (2023)).
In view of these challenges and opportunities, we propose a novel framework K-M3AID (Multi-Level Multimodal Alignment with Knowledge-Guided Instance-Wise Discrimination), aiming at achieving reliable analog retrieval, candidate ranking, and peak assignment in the interpretation of NMR spectra (see Figure  1.b). The overview of our K-M3AID framework is a dual-coordinated contrastive learning architecture, which contains three key components: graph-level alignment module, node-level alignment module, and communication channel. Graph-level alignment module establishes the correspondences of molecules with their individual ${}^{13}$ NMR spectra. Given that each unique molecule gives rise to a distinct spectral signature, this module is equipped with a straightforward cross-entropy loss designed for effective contrastive learning. Node-level alignment module aligns each C atom within the molecules with their signal peaks on the spectrum. In contrast to the diverse and distinctive molecular spectral signatures, many atoms exhibit chemical symmetric and magnetic equivalence within the same molecule, corresponding to the same peaks. Meanwhile, atoms with different local surroundings can still presents significant similarity on the spectrum, which introduces heightened level of complexity. In view of these complex scenarios, we come up with knowledge-guided instance wise discrimination based contrastive learning in graph-level alignment module (see Figure 4). The communication channel dynamically adjusts the flow of gradients between the node encoder and the graph encoder from two modules during the training process.
In summary, our contribution comprises three major aspects: Conceptually: we integrate two architectural levels alignment, graph- and node-levels, into the K-M3AID framework. This integration facilitates rapid adaptation, significantly boosting the efficiency of learning for multimodal zero-shot tasks. Methodologically: We present knowledge-guided instance wise discrimination for cross-modal contrastive learning, which take advantage of continuous and domain-specific features with natural ordering. To the best of our knowledge, this is the first to demonstrate knowledge-guided instance wise discrimination based cross-modal contrastive learning, which transforms a discrete comparison into a continuous one. Empirically: We demonstrate the effectiveness of K-M3AID through its successful application to multiple zero-shot tasks, including molecular retrieval, isomer recognition, and peak assignment.
Figure 1: a. Demands for interpreting NMR spectra in real-world scenarios: molecular retrieval, candidate ranking, and peak assignment; b. Zero-shot applications of the K-M3AID model: molecular retrieval, isomer recognition, and peak assignment; c. The framework of K-M3AID model: the molecular alignment module is responsible for optimizing the the correspondence between modalities at the molecular level, the atomic alignment module focus on the fine-tuning of atomic positioning on the spectrum, and the communication channel dynamically adjusts the flow of gradients between node encoder and graph encoder during the training process. S for spectrum embedding, G for graph embedding, P for peak embedding and N for node embedding.
2 Preliminaries
Multimodal Alignment: Multimodal alignment, as defined in the literature (Baltrusaitis et al. (2017)), involves establishing relationships and correspondences among sub-components of instances from two or more modalities. A typical example is identifying specific regions in an image that correspond to words or phrases in a given caption (Karpathy & Fei-Fei (2015)). This approach offers numerous benefits, including enhanced data interpretation, heightened accuracy and robustness, overcoming limitations of single-modal systems, and better addressing real-world complexity (Baltrusaitis et al. (2017), Summaira et al. (2021), Akkus et al. (2023)). CLIP (Contrastive Language-Image Pretraining) (Radford et al. (2021); Li et al. (2021)) is one of the most widely adopted frameworks for multimodal alignment. As highlighted in the introduction, molecular information originates from diverse sources such as molecule graphs and NMR spectroscopy. To leverage effective alignments of this multifaceted information across different modalities, we adopt CLIP framework with graph neural networks (GNN) (Xu et al. (2018), Wu et al. (2022)) to encoder molecular information and neural network encoders (Serra et al. (2018)) to encapsulate NMR information.
Meta-Learning: Meta-learning is defined as learning how to learn across tasks (Vilalta & Drissi (2002)). More specifically, it leverages skills previously acquired from related tasks to the current one (Lake et al. (2017). With more skills learned, learning new skills becomes easier, requiring fewer examples and less trial-and-error (Vanschoren (2018); Finn et al. (2017)). A meta-learner is trained on a diverse set of object recognition tasks. During this training, it learns common features, patterns, and strategies for recognizing objects. Once trained, when presented with a new, previously unseen object category, the meta-learner can rapidly adapt and achieve high recognition accuracy, leveraging the knowledge it acquired from the diverse training tasks to perform in this novel recognition task (Finn et al. (2017). A profound understanding of atom properties allows us to extend our vision to previously unseen molecules, aligning with the principles of meta-learning in artificial intelligence.
Contrastive Learning: Contrastive learning focus on learning how to distinguish between similar and dissimilar items (Le-Khac et al. (2020b); Jaiswal et al. (2021); Liu et al. (2021)). A fundamental aspect of this process involves instance-wise discrimination (Wu et al. (2018)). Models incorporating instance-wise discrimination not only foster an understanding of the inherent data structure but also enhance generalization capabilities. This is attributed to the contrastive learning approach, which prioritizes distinguishing between instances rather than memorizing specific labeled examples. In chemistry, each molecule/atom is treated as a distinct instance. The learning algorithm focuses on distinguishing each molecule/atom based on its context.
3 Experiments
To thoroughly evaluate the performance of K-M3AID, we compare it with other baselines through various zero-shot downstream tasks, such as molecular retrieval, isomer recognition and peak assignment.
3.1 Experimental Setting
3.1.1 Pre-training
Dataset: We use ${}^{13}$ NMR spectra of about 20,000 molecules sourced from nmrshiftdb2 (Steinbeck et al. (2003)), a public access database that contains NMR spectra of organic molecules. In the collected dataset, molecule are aligned with their respective ${}^{13}$ NMR spectra, and atomic alignments with peaks are also included. Notably, the dataset contains 12,771 molecules with fewer than 10 carbon atoms, 7,043 molecules featuring carbon atom counts ranging from 10 to 20, and 1,138 molecules incorporating more than 20 carbon atoms. The quality of the dataset was further validated by experienced organic chemists. We randomly sample 80% of the molecules for training and the rest for evaluation.
Training: We concurrently leverage both graph- and node-level alignment tasks in the pre-training of K-M3AID. Graph-level alignment focuses on aligning molecules with their spectra, accompanied by the utilization of cross-entropy loss for contrastive purposes. On the other hand, node-level alignment entails aligning atoms with their corresponding peaks, implemented through knowledge-guided instance-wise discrimination to achieve contrastive loss. A diverse set of molecular features is employed for training, including atomic number (node feature), chiral tags (node feature), hybridization (node feature), bond types (edge feature), and bond direction (edge feature). The spectral features are derived from peak intensity, peak position (chemical shift measured in ppm), and peak type.
3.1.2 Zero-Shot Molecular Retrieval
Dataset: We randomly collected 1 million molecules from PubChem (Kim et al. (2023)) to form a molecular reference library. Subsequently, we carefully selected 1000 spectra, ensuring that they had not appeared in the training dataset, from an external dataset to serve as query spectra. Following this, the corresponding molecules associated with these 1000 spectra were added into the existing reference library.
Evaluation: We perform molecular retrieval using each of the selected spectra to determine if the correct corresponding molecular entity can be retrieved from the reference library. The model’s performance is assessed at top-1% accuracy, as well as at accuracy of 5%, 10%, and 25%.
3.1.3 Zero-Shot Isomer Recognition
Dataset: We categorize isomers from the validation dataset to guarantee their absence from the training dataset. To assess effectiveness, we perform isomer recognition on each isomer group containing at least 10 molecules. Within the same group, isomers may be structural or spatial isomers of each other. Structural isomers refer to molecules with the same molecular formula but different structural arrangements of atoms, resulting in distinct chemical structures. On the other hand, spatial isomers, also known as stereoisomers, have the same molecular formula and arrangement of atoms but differ in the spatial orientation of their atoms in three-dimensional space, leading to different stereoisomeric forms. (For an elucidation of an isomer group and in-depth insights into isomers with NMR, please refer to the details provided in Appendix D.)
Evaluation: We conduct isomer recognition for each isomer group, aiming to assess the correct alignment of each spectrum with its respective molecule within each isomer group.
3.1.4 Zero-Shot Peak Assignment
Dataset: The dataset utilized for evaluating the overall performance of K-M3AID on zero-shot peak assignment is the validation dataset from the pre-training phase. In order to highlight the capabilities of K-M3AID in zero-shot peak assignment, the case studies include complex natural products featuring multiple fused rings, stereogenic (chiral) centers, and symmetric structures.
Evaluation: We conduct peak assignment within each molecule, aiming to assess the accurate alignment of each atom with its corresponding peak on the spectrum. It’s important to note that this alignment process is confined to each individual molecule and not across different molecules.
3.1.5 Baselines
In contrast to communicative mechanism of K-M3AID, one of the baselines is established without the utilization of a communication channel. Additionally, we explore two alternative baselines where the knowledge-guided instance-wise discrimination in the node-level alignment module is replaced with either strong-pair-based (SP-ID) or weak-pair-based (WP-ID) instance-wise discrimination. SP-ID enforces a precise match in node-level (atom-peak) alignment, ensuring that only correct pairs established during the training process are considered. On the other hand, WP-ID broadens the matching criteria, allowing for multiple matches within a specified threshold set for the distance of their corresponding parts per million (ppm, referenced in Section  6.4). The mathematical definitions of strong pair and weak pair are as follows:
[…]
Here, $A_{i}\in A$ represents atoms, $P_{j}\in P$ represents peaks, and $th$ is the abbreviation for the threshold. Additionally, $ppm_{i}$ denotes the parts per million for the $i^{th}$ Carbon Atom, and $ppm_{j}$ represents the corresponding parts per million for the $j^{th}$ peak.
3.2 Results
3.2.1 Validation Performance
The K-M3AID model demonstrates an impressive validation accuracy of 95.5% for molecule-spectrum alignments within the graph-level alignment module. In direct comparison, K-M3AID surpasses alternative models such as SP-ID, WP-ID, and the model without a communication mechanism, exhibiting superior performance by a margin ranging from about 1% to 6% in the graph-level alignment module (refer to Table 1). Notably, SP-ID outperforms WP-ID significantly, and as the matching criteria threshold widens, the performance of the latter degrades.
In the context of peak-atom alignment within the node-level alignment module, K-M3AID and the model without a communication mechanism exhibit comparable accuracies (refer to Table 1). However, K-M3AID showcases slightly better stability across 5-fold cross-validation. Moreover, K-M3AID demonstrates superiority in peak-atom alignment when compared to SP-ID and WP-ID. This superiority may arise from the inherent limitations of both strong and weak pair definitions, which is failing to precisely calibrate the diverse relationships among the elements. This finding is further supported by the significant decreases in the accuracy of peak-atom alignment as the threshold of weak pair increases.
Table 1: Batch-wise validation accuracy (%) of K-M3AID and baselines with $epochs=200$.
Alignment SP-ID WP-ID(th=1) WP-ID(th=5) WP-ID(th=10) No Comm. K-M3AID
Graph-Level 93.5$\pm$0.6 91.3$\pm$0.8 90.3$\pm$0.6 88.4$\pm$1.4 94.6$\pm$0.4 95.5$\pm$0.4
Node-Level 89.3$\pm$0.4 83.7$\pm$0.6 79.8$\pm$0.5 66.1$\pm$2.5 90.4$\pm$0.2 90.3$\pm$0.1
3.2.2 Performance on Zero-Shot Molecular Retrieval
We conduct a systematic evaluation of the effectiveness of our K-M3AID model, comparing it with baseline models in the zero-shot molecular retrieval task across datasets of varying magnitudes. Detailed results are presented in Table 2. The K-M3AID model consistently attains an impressive top-1 accuracy of approximately 95.8% in molecular retrieval when the molecular reference library comprises 100 entries. This performance surpasses that of alternative mechanisms such as SP-ID (95.3%), WP-ID (92.9%), and No Communication (94.8%). As the molecular reference library expands to 1000 entries, the K-M3AID model exhibits notable superiority, achieving accuracy levels of 80.4%, 1.8%, 8.7%, and 2.8% higher than SP-ID, WP-ID, and No Communication mechanisms, respectively. The advantage of K-M3AID becomes even more pronounced when the library size reaches 10,000 entries. In this scenario, K-M3AID yields 46.3% at top-1 accuracy, showcasing advancements of 10.5%, 13.6%, and 6.2% over SP-ID, WP-ID, No Communication mechanisms, respectively. Even with larger molecular reference libraries, such as 100,000 and 1,000,000 entries, K-M3AID consistently outshines SP-ID, WP-ID, and No Communication mechanisms. These compelling results distinguish the K-M3AID model as an exceptional choice in scenarios demanding robust performance in molecular retrieval tasks.
Table 2: Zero-shot molecular retrieval at top 1(%) accuracy across datasets of varying sizes. (For more statistics, such as top 5(%), top 10(%) and top 25(%), please refer to Appendix Table  C.1.)
Method $10^{2}$ $10^{3}$ $10^{4}$ $10^{5}$ $10^{6}$
K-M3AID 95.8$\pm$1.0 80.4$\pm$3.9 46.3$\pm$1.2 18.0$\pm$0.8 5.8$\pm$1.7
SP-ID 95.3$\pm$0.8 78.6$\pm$2.7 35.8$\pm$3.8 12.9$\pm$1.6 3.4$\pm$0.9
WP-ID (th = 1) 92.9$\pm$0.6 71.7$\pm$1.0 32.7$\pm$1.3 10.7$\pm$0.5 3.6$\pm$0.7
No Comm. 94.8$\pm$1.2 77.6$\pm$1.4 40.1$\pm$1.2 14.4$\pm$0.9 4.1$\pm$1.1
3.2.3 Performance on Zero-Shot Isomer Recognition
K-M3AID stands out prominently when compared to SP-ID, WP-ID, and no communication approaches in the task of zero-shot isomer recognition, achieving an exceptional 100% accuracy across given groups of isomers (refer to Table 3). These empirical observations underscore the advantages of K-M3AID in the context of isomer recognition. The superiority of K-M3AID over the no communication baseline demonstrates the positive impact of node-level alignment on graph-level alignment, emphasizing the potency of meta-learning.
Table 3: Zero-shot isomer recognition accuracy (%) of K-M3AID and baselines.
Formula #Isomers SP-ID WP-ID (th=1) No Comm. K-M3AID
$\text{C}_{4}\text{H}_{6}\text{O}$ 15 86.7 86.7 86.7 100.0
$\text{C}_{9}\text{H}_{9}\text{N}$ 15 86.7 80.0 100.0 100.0
$\text{C}_{7}\text{H}_{11}\text{NO}_{3}$ 14 78.6 85.7 85.7 100.0
$\text{C}_{6}\text{H}_{13}\text{NO}$ 23 91.3 91.3 100.0 100.0
$\text{C}_{8}\text{H}_{7}\text{NO}_{4}$ 13 92.3 84.6 92.3 100.0
$\text{C}_{15}\text{H}_{24}\text{O}$ 16 93.8 93.8 100.0 100.0
$\text{C}_{11}\text{H}_{14}$ 10 90.0 80.0 70.0 100.0
$\text{C}_{7}\text{H}_{15}\text{NO}$ 14 85.7 85.7 100.0 100.0
$\text{C}_{10}\text{H}_{16}\text{O}_{2}$ 26 92.3 84.6 100.0 100.0
$\text{C}_{8}\text{H}_{15}\text{N}$ 11 81.8 90.9 100.0 100.0
3.2.4 Performance on Zero-Shot Peak Assignment
K-M3AID model demonstrates a validation accuracy surpassing 90% for peak assignment (peak-atom alignment) within the node-level alignment module after 200 epochs (see Figure 2.A). Notably, the model achieves a 100% accuracy rate in 74.1% of molecules containing fewer than 10 carbon atoms (see Figure 2.B). For molecules with carbon atom counts ranging from 10 to 20, the model attains 100% accuracy in 37.2% of cases (see Figure 2.C). Furthermore, it achieves an accuracy exceeding 80% in more than 50% of cases pertaining to molecules containing more than 20 carbon atoms (see Figure 2.D). Additionally, to further illustrate the power of K-M3AID on peak assignment, we present two complex natural product molecules featuring multiple rings (4 and 4, respectively) and stereogenic (chiral) centers (6 and 8, respectively) in Figure E.1.
Figure 2: The statistics of zero-shot peak assignment.
K-M3AID demonstrates superior performance in peak assignment compared to SP-ID and WP-ID. Our case studies reveal that the limitations of SP-ID and WP-ID become particularly pronounced in two scenarios: 1) when local contexts of specific atoms exhibit a high degree of similarity; 2) when certain atoms display symmetric mapping within the same molecule. In the former scenario, exemplified by molecular A in Figure  3, atom 0 and atom 4 are secondary carbons (attaching to 2 carbons and 2 hydrogens), nearly symmetric on the same 5-member ring, corresponding to the peak position measured in ppm of 27.0 and 29.8, respectively. (For the definition of ppm, please refer to Section 6.4.) The similar local content of these two atoms fools SP-ID and WP-ID. Meanwhile, atom 1 and atom 3 are tertiary carbons (attaching to 3 carbons and 1 hydrogen), nearly symmetric on the same 5-member ring, corresponding to the peak position measured in ppm of 54.5 and 44.1, respectively. Only WP-ID fails to distinguish and align them. In the later scenario, exemplified by molecular B in Figure  3, there exist instances one-to-one and one-to-many for atomic-level alignment within the molecular configuration. Both SP-ID and WP-ID method misaligns certain atoms with other atoms with small ppm differences (less than 3 ppm in this case), rather than aligning them with themselves or their symmetric counterparts. In contrast, K-M3AID approach excels in both scenario by discerning each one of the atoms, which is attributed to the full utilization of ppm difference distance learning. (For additional cases, please refer to Appendix Figure E.2)
Figure 3: Case study of peak assignment. Yellow cells in PPM difference represent the ground truth alignment, and red cross represents the wrong alignment. For the definition of ppm, please refer to 6.4. For additional cases, please refer to Appendix Figure E.2
4 Related Work
Multimodal Instance-Wise Discrimination: As mentioned in the prelimtaries, instance discrimination (Le-Khac et al. (2020a); Zolfaghari et al. (2021); Morgado et al. (2021); Liu et al. (2023)), an important part of contrastive learning, distinguishes individual instances without explicit class labels. Moving into multimodal contrastive learning, it can be categorized into two general approaches: strong-pair-based (van den Oord et al. (2019); Jaiswal et al. (2021); Liu et al. (2023)) and weak-pair-based (Salakhutdinov & Hinton (2007); Frosst et al. (2019); Liang et al. (2021)) instance-wise discrimination. The strong-pair-based such as NCE method enforces a precise one-to-one correspondence for real samples with artificially generated noise samples. An example of positive pair can be a noise-added picture of zebra with the text description of zebra. Instead of one-to-one correspondences, weak-pair-based approach relaxes the positive pairs to more boarder semantic correspondences. An example of positive pair can be a picture of zebra with the text description of horse, but not with the text description of tiger.
Multimodal Meta-Alignment: Within the scope of multimodal alignment, multimodal meta-alignment is a novel method to align representation spaces using paired cross-modal data with different similarity while ensuring quick generalization to new tasks across different modalities (Liang et al. (2021)). This approach can be observed at different levels, including the intermediate and fundamental (irreducible) element level. Examples of this method at the intermediate level can be seen in research on Cross-Modal Generalization (Chen et al., 2017; Li et al., 2020; Liang et al., 2021; Zhang et al., 2021) and Livestreaming Product Recognition (Yang, 2023). While these studies showcase how multimodal meta-alignment operates at a broad objective level, the application of multimodal meta-alignment at the most fundamental element level remains underexplored in current research.
5 Conclusion and Future Work
In this paper, we introduced K-M3AID (Knowledge-Guided Multi-Level Multimodal Alignment with Instance-Wise Discrimination) framework, incorporating graph-level alignment and node-level alignment. Its effectiveness was demonstrated through multiple zero-shot tasks: molecular retrieval, isomer recognition and peak assignment. The significance of knowledge-guided instance-wise discrimination is underscored through various metrics and case studies. Moreover, the findings from molecular retrieval and isomer recognition underscored the favorable influence of the node-level alignment on the graph-level alignment. It emphasizes the successful integration of meta-learning within our hierarchical alignment framework. While our framework achieves an atomic-level alignment overall accuracy of 100% for 55% of cases, it drops significantly to 9.8% when handling molecules with more than 20 carbon atoms. Presently, our graph encoder operates on 2D molecular graphs with basic node and edge features. This implementation potentially constrains its ability to generate precise node embeddings for distinguishing atoms in highly complex scenarios. Future developments could benefit from incorporating a 3D-based graph, holding substantial potential to enhance performance in such complex situations.
6 Our Method
In this section, we firstly present the architecture of K-M3AID framework, an end-to-end system designed for multi-level multimodal alignment. Then, we introduce the constrastive learning loss in K-M3AID along with the principles of Knowledge-Guided Instance-Wise Discrimination.
6.1 Architecture
The K-M3AID framework is a dual-CLIP architecture (see Figure 1), which consists of three critical components: graph-level alignment module, node-level alignment module and communication channel. The graph-level alignment module adapts a gradient-asymmetric CLIP mechanism. While two unimodal encoders work in conjunction, only the from-scratch graph encoder (GIN, Xu et al. (2018)) undergoes dynamic training throughout the process, the pre-trained spectrum encoder (Yang et al. (2021)) remains fixed. Both encoders are complemented by dedicated projection layers, which facilitate the mapping of embeddings into a joint space. The node-level alignment module adapts a gradient-symmetric CLIP mechanism. It is equipped with two from-scratch unimodal encoders, node encoder and peak encoder, as well as their dedicated projection layers. The graph encoder in graph-level alignment module shares part of the weights with node encoder in node-level alignment module, serving as the communication channel.
6.2 Contrastive Learning Loss
The synergy between these two modules is pivotal, collectively contributing to our loss function, expressed as
$L=CL_{graph}+CL_{node},$ (3)
where $CL_{graph}$ represents the contrastive learning loss in the graph-level alignment module by Equation 5, and $CL_{node}$ represents the contrastive learning loss in the node-level alignment module by Equation 15.
Let $i$ denote the $i^{th}$ instance, and $j$ denote the $j^{th}$ instance. Then $x_{i}$ denotes the raw input in modality A for the $i^{th}$ instance and $y_{j}$ denotes the raw input in modality B for the $j^{th}$ instance. Suppose $f_{x}\left(\cdot\right)$ represent the encoding function for modality A, and $f_{y}\left(\cdot\right)$ denote the encoding function for modality B. In graph-level alignment module, these two encoding functions, should map $x_{i}$ and $y_{j}$ to a proximate location in the joint embedding (inter-modality) if $i=j$.
$CL_{graph}(i)$ $=-\log\frac{e^{\delta(x_{i},y_{i})}}{\displaystyle\sum_{1\leq j\leq N}e^{\delta(x_{i},y_{j})}}$ (4)
$=-\text{log}(\text{softmax}(\delta(x_{i},y_{i}))$ (5)
Where $\delta(x_{i},y_{j})=\left(f_{x}(x_{i})^{T}\cdot f_{y}(y_{j})\right)$, $N$ is the total number of instances from the current batch.
Thus, the total $CL_{graph}$ is expressed as following:
$CL_{graph}=\frac{1}{N}\displaystyle\sum_{1\leq i\leq N}CL_{graph}(i)$ (6)
This design for the loss aims to match the same instance cross different modalities.
6.3 Knowledge-Guided Instance-Wise Discrimination Contrastive Learning
Figure 4: Knowledge span mechanism. $K_{i}$ and $K_{j}$ represents the corresponding knowledge span label for $i^{th}$ and $j^{th}$ items.
Knowledge Span, in which we define as some continuous and domain-specific feature exhibiting natural ordering and offering guidance, can potentially offer valuable insights into the contrastive learning process. As such, we propose a novel and general approach to contrastive learning called Knowledge-Guided Instance-Wise Discrimination (see Figure  4). This approach expands the scope of contrastive learning from confined comparisons (pre-determined negative and positive pairs) to unrestricted comparisons (no need for pre-determination). This extension removes the necessity of explicitly defining such pairs, thus mitigating the potential introduction of human bias.
Suppose $\mathcal{M}$ is the set of instances. $\mathcal{A}\subset\mathbb{R}^{d_{1}}$ is the set of tunable instances’ embeddings in modality A, $\mathcal{B}\subset\mathbb{R}^{d_{1}}$ is the set of tuable instances’ embeddings in modality B, and $\mathcal{K}\subset\mathbb{R}^{d_{2}}$ is the corresponding fixed knowledge span label that can guide the relative distance learning between components in $\mathcal{A}$ and $\mathcal{B}$. Thus, the size of $\mathcal{A}$, $\mathcal{B}$, $\mathcal{K}$ are $|\mathcal{M}|$, respectively.
Let $\mathcal{A}_{i}$ be the $i^{th}$ instance embedding of $\mathcal{A}$, and $\mathcal{B}_{j}$ be the $j^{th}$ instance embedding of $\mathcal{B}$. We define the distance function between $\mathcal{A}_{i}$ and $\mathcal{B}_{j}$ as $d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})=\mathcal{A}_{i}\cdot\mathcal{B}_{j}\rightarrow\mathbb{R}^{+}$, and calibration function $d(\mathcal{K}_{i},\mathcal{K}_{j})\rightarrow\mathbb{R}^{+}$ with a monotonic property and constraint $\sum_{j=1}^{|\mathcal{M}|}d(\mathcal{K}_{i},\mathcal{K}_{j})=1$, in which $\mathcal{K}_{i}$ and $\mathcal{K}_{j}$ serve as the designated Knowledge Span Label. We introduce the Knowledge Span Guided Loss (KSGL) as follows:
$KSGL(i)$ $=-\displaystyle\sum_{\begin{subarray}{c}1\leq j\leq|\mathcal{M}|\end{subarray}}d(\mathcal{K}_{i},\mathcal{K}_{j})\log\frac{e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})}}{\displaystyle\sum_{\begin{subarray}{c}1\leq k\leq|\mathcal{M}|\end{subarray}}e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{k})}}$ (9)
$=-\displaystyle\sum_{\begin{subarray}{c}1\leq j\leq|\mathcal{M}|\end{subarray}}d(\mathcal{K}_{i},\mathcal{K}_{j})\log(\text{softmax}(d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})))$ (11)
In particular, when it reaches ideal optimum, $d(\mathcal{K}_{i},\mathcal{K}_{j})$ and $d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})$ reaches the following relation:
$d(\mathcal{K}_{i},\mathcal{K}_{j})=\text{softmax}(d_{E}(\mathcal{A}_{i},\mathcal{B}_{j}))$ (12)
For detail proof, please refer to Appendix A. As a result, the corresponding $CL_{instance}$ is expressed as following:
$CL_{instance}$ $=\frac{1}{|\mathcal{M}|}\displaystyle\sum_{1\leq i\leq|\mathcal{M}|}KSGL(i)$ (13)
6.4 Chosen Knowledge Span-ppm
${}^{13}$ NMR uncovers molecular structures by providing the chemical environments of carbon atoms and their magnetic responses to external fields, and quantifies these features in parts per million (ppm) relative to a reference compound like tetramethylsilane (TMS), simplifying comparisons across experiments. Thus, continuous peak positions, measured in ppm, can serve as a robust knowledge span to facilitate instance-wise discrimination for this contrastive learning task.
For node-level alignment module, $\mathcal{A}$ is the set of node embeddings for Carbon atoms in molecular graph modality and $\mathcal{B}$ is the set of peak embeddings for respective Carbon atoms in NMR modality, $\mathcal{K}$ is the set of ppm value for each corresponding Carbon atom in $\mathcal{A}$ and $\mathcal{B}$. Suppose $ppm_{i}$ is the ppm for the $i^{th}$ Carbon Atom and $ppm_{j}$ is the corresponding ppm for $j^{th}$ peak. $d(\cdot,\cdot)$ is then defined as following:
$d(\mathcal{K}_{i},\mathcal{K}_{j})=d(ppm_{i},ppm_{j})$ $=softmax(\frac{\tau_{2}}{|ppm_{i}-ppm_{j}|+\tau_{1}})$ (14)
where $\tau_{1}$ and $\tau_{2}$ are temperature hyper-parameter. For further discussion of selection about $\tau_{1}$ and $\tau_{2}$, please refer to Appendix B.2. Then, the final form of contrastive loss for node-level alignment according to Equation 11 and Equation 13 is as following:
$CL_{node}=-\frac{1}{|\mathcal{M}|}\sum_{\begin{subarray}{c}1\leq j\leq|\mathcal{M}|\end{subarray}}d(ppm_{i},ppm_{j})\log\frac{e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})}}{\sum_{\begin{subarray}{c}1\leq k\leq|\mathcal{M}|\end{subarray}}e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{k})}}$ (15)
Appendix
Appendix A Revisiting Knowledge Span Guided Loss
Theorem 1 (Knowledge Span Guided Loss).
Suppose $\mathcal{M}$ is the set of instances. $\mathcal{A}\subset\mathbb{R}^{d_{1}}$ is the set of tunable instances’ embeddings in modality A, $\mathcal{B}\subset\mathbb{R}^{d_{1}}$ is the set of tuable instances’ embeddings in modality B, and $\mathcal{K}\subset\mathbb{R}^{d_{2}}$ is the corresponding fixed knowledge span label that can guide the relative distance learning between components in $\mathcal{A}$ and $\mathcal{B}$. Thus, the size of $\mathcal{A}$, $\mathcal{B}$, $\mathcal{K}$ are $|\mathcal{M}|$, respectively.
Let $\mathcal{A}_{i}$ be the $i^{th}$ instance embedding of $\mathcal{A}$, and $\mathcal{B}_{j}$ be the $j^{th}$ instance embedding of $\mathcal{B}$. We define the distance function between $\mathcal{A}_{i}$ and $\mathcal{B}_{j}$ as $d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})=\mathcal{A}_{i}\cdot\mathcal{B}_{j}\rightarrow\mathbb{R}^{+}$, and calibration function $d(\mathcal{K}_{i},\mathcal{K}_{j})\rightarrow\mathbb{R}^{+}$ with a monotonic property and constraint $\sum_{j=1}^{|\mathcal{M}|}d(\mathcal{K}_{i},\mathcal{K}_{j})=1$, in which $\mathcal{K}_{i}$ and $\mathcal{K}_{j}$ serve as the designated Knowledge Span Label. We introduce the Knowledge Span Guided Loss (KSGL) as follows:
$KSGL(i)$ $=-\displaystyle\sum_{\begin{subarray}{c}1\leq j\leq|\mathcal{M}|\end{subarray}}d(\mathcal{K}_{i},\mathcal{K}_{j})\log\frac{e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})}}{\displaystyle\sum_{\begin{subarray}{c}1\leq k\leq|\mathcal{M}|\end{subarray}}e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{k})}}$ (A.3)
[…]
Proof.
In order to optimize the loss $KSGL(i)$, we need to set the following partial derivative to be 0 for each $d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})$ with $1\leq j\leq|\mathcal{M}|$. Here are the detail process:
$\frac{\partial KSGL(i)}{\partial{d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})}}$ $=\frac{\partial}{\partial d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})}\underbrace{\left(-d(\mathcal{K}_{i},\mathcal{K}_{j})\log\frac{e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})}}{e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})}+\sum_{\begin{subarray}{c}k\neq j\end{subarray}}e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{k})}}\right)}_{\text{When the numerator includes }e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})}}$
$\quad+\frac{\partial}{\partial d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})}\underbrace{\left(\sum_{\begin{subarray}{c}k\neq j\end{subarray}}-d(\mathcal{K}_{i},\mathcal{K}_{k})\log\frac{e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{k})}}{e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})}+\sum_{\begin{subarray}{c}k\neq j\end{subarray}}e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{k})}}\right)}_{\text{When the numerator does not include }e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})}}$
$=-(d(\mathcal{K}_{i},\mathcal{K}_{j})-d(\mathcal{K}_{i},\mathcal{K}_{j})\cdot\text{softmax}(d_{E}(\mathcal{A}_{i},\mathcal{B}_{j}))$
$\quad-\sum_{\begin{subarray}{c}k\neq j\end{subarray}}d(\mathcal{K}_{i},\mathcal{K}_{k})\cdot\text{softmax}(d_{E}(\mathcal{A}_{i},\mathcal{B}_{j}))$
$=-\left(d(\mathcal{K}_{i},\mathcal{K}_{j})-(d(\mathcal{K}_{i},\mathcal{K}_{j})+\sum_{\begin{subarray}{c}k\neq j\end{subarray}}d(\mathcal{K}_{i},\mathcal{K}_{k}))\cdot\text{softmax}(d_{E}(\mathcal{A}_{i},\mathcal{B}_{j}))\right)$
Since $\sum_{l=1}^{|\mathcal{M}|}d(\mathcal{K}_{i},\mathcal{K}_{l})$ = 1, we can further simplify it as
$\frac{\partial KSGL(i)}{\partial{d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})}}=-(d(\mathcal{K}_{i},\mathcal{K}_{j})-\text{softmax}(d_{E}(\mathcal{A}_{i},\mathcal{B}_{j}))$
In order to optimize, we need to set the respective partial derivative to be 0:
$\frac{\partial KSGL(i)}{\partial{d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})}}=-(d(\mathcal{K}_{i},\mathcal{K}_{j})-\text{softmax}(d_{E}(\mathcal{A}_{i},\mathcal{B}_{j}))=0$
In addition, the corresponding second partial derivative denoted as $\frac{\partial KSGL(i)}{\partial{d_{E}^{2}(\mathcal{A}_{i},\mathcal{B}_{j})}}$ manifests as follows:
$\frac{\partial KSGL(i)}{\partial{d_{E}^{2}(\mathcal{A}_{i},\mathcal{B}_{j})}}=\text{softmax}(d_{E}(\mathcal{A}_{i},\mathcal{B}_{j}))(1-\text{softmax}(d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})))$
As $\text{softmax}(d_{E}(\mathcal{A}_{i},\mathcal{B}_{j}))$ takes values within the open interval (0,1), it follows that $\frac{\partial KSGL(i)}{\partial{d_{E}^{2}(\mathcal{A}_{i},\mathcal{B}_{j})}}$ is always positive. Consequently, the pinnacle of optimization emerges as a global minimum.
Furthermore, when it comes to optimum:
$d(\mathcal{K}_{i},\mathcal{K}_{j})$ $=\text{softmax}(d_{E}(\mathcal{A}_{i},\mathcal{B}_{j}))$
$d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})$ $=\log(d(\mathcal{K}_{i},\mathcal{K}_{j}))+\log\left(\sum_{\begin{subarray}{c}1\leq l\leq|\mathcal{M}|\end{subarray}}e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{l})}\right)$
It is easy to show that when it reaches optimum, $d_{E}(A_{i},B_{j})$ is consistent with Knowledge Span Guidance $d(\mathcal{K}_{i},\mathcal{K}_{j})$. Without loss of generosity, suppose $d(\mathcal{K}_{i},\mathcal{K}_{j})>d(\mathcal{K}_{i},\mathcal{K}_{j^{\prime}})$ :
$d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})-d_{E}(\mathcal{A}_{i},\mathcal{B}_{j^{\prime}})$ $=\log(d(\mathcal{K}_{i},\mathcal{K}_{j}))+\log\left(\sum_{\begin{subarray}{c}1\leq l\leq|\mathcal{M}|\end{subarray}}e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{l})}\right)$
$\quad-\left(\log(d(\mathcal{K}_{i},\mathcal{K}_{j^{\prime}}))+\log\left(\sum_{\begin{subarray}{c}1\leq l\leq|\mathcal{M}|\end{subarray}}e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{l})}\right)\right)$
$=\log(d(\mathcal{K}_{i},\mathcal{K}_{j}))-\log(d(\mathcal{K}_{i},\mathcal{K}_{j^{\prime}}))$
$=\log\left(\frac{d(\mathcal{K}_{i},\mathcal{K}_{j})}{d(\mathcal{K}_{i},\mathcal{K}_{j^{\prime}})}\right)>0$
Appendix B Further ablation study about parameter choices
B.1 Ablation study about the choice of GIN structure and projection.
We choose GIN(Xu et al. (2018)) as our graph encoder. By Table B.1, ”GIN Depth” signifies the number of layers in the GIN, ”GIN Embedding Dim” denotes the dimensionality of the embeddings generated by the GIN model, and ”Projection Dim” indicates the resulting dimensionality after transforming the GIN-produced embeddings. In particular, the best performance is observed when the GIN model has 5 layers, GIN Embedding Dim is 128, and projection Dim is 512.
Table B.1: GIN structure and projection ablation study
GIN Depth GIN Embedding Dim Projection Dim Validation accuracy (%)
3 128 128 86.6
3 256 128 86.8
3 512 128 86.3
5 128 128 89.4
5 256 128 89.6
5 512 128 89.3
3 128 256 86.6
3 256 256 86.8
3 512 256 86.3
5 128 256 89.4
5 256 256 89.6
5 512 256 89.3
3 128 512 86.6
3 256 512 86.5
3 512 512 86.2
5 32 512 84.0
5 64 512 87.5
5 128 512 90.0
5 256 512 89.4
5 512 512 88.9
B.2 Ablation Study on the Choice of $\tau_{1}$ and $\tau_{2}$
We also conducted a further ablation study exploring different combinations of $\tau_{1}$ and $\tau_{2}$ as shown in Table B.2. For this analysis, we fixed the GIN depth at 5, set the GIN embedding dimensionality to 128, and maintained a projection dimension of 512. We observe that the best performance is achieved when $\tau_{1}$ = $10^{-5}$ and $\tau_{2}$ = $10^{1}$.
Table B.2: Ablation study about $tau_{1}$ and $tau_{2}$. We have 5 layers and 128 dimension as the final representation.
$\tau_{1}$ $\tau_{2}$ Molecular Alignment Accuracy (%) Atom Alignment Accuracy (%)
$10^{-1}$ $10^{1}$ 94.9 89.6
$10^{-1}$ $10^{2}$ 95.2 89.8
$10^{-1}$ $10^{3}$ 95.6 89.6
$10^{-1}$ $10^{4}$ 95.1 88.9
$10^{-1}$ $10^{5}$ 95.0 89.3
$10^{-2}$ $10^{1}$ 95.5 89.8
$10^{-2}$ $10^{2}$ 94.8 89.8
$10^{-2}$ $10^{3}$ 95.4 88.8
$10^{-2}$ $10^{4}$ 94.8 87.2
$10^{-2}$ $10^{5}$ 95.1 89.4
$10^{-3}$ $10^{1}$ 95.0 89.2
$10^{-3}$ $10^{2}$ 95.1 89.1
$10^{-3}$ $10^{3}$ 95.2 89.0
$10^{-3}$ $10^{4}$ 95.3 89.7
$10^{-3}$ $10^{5}$ 95.0 89.4
$10^{-4}$ $10^{1}$ 95.0 89.8
$10^{-4}$ $10^{2}$ 95.1 89.7
$10^{-4}$ $10^{3}$ 95.0 89.8
$10^{-4}$ $10^{4}$ 95.3 89.5
$10^{-4}$ $10^{5}$ 95.1 88.4
$10^{-5}$ $10^{1}$ 95.4 90.0
$10^{-5}$ $10^{2}$ 95.0 89.5
$10^{-5}$ $10^{3}$ 95.8 89.6
$10^{-5}$ $10^{4}$ 95.2 89.7
$10^{-5}$ $10^{5}$ 95.0 89.7
Appendix C Additional Results on Molecular Retrieval
Table C.1: Zero-shot molecular retrieval top 5, 10, 25 accuracy (%) with K-M3AID and baselines
Method Accuracy $10^{2}$ $10^{3}$ $10^{4}$ $10^{5}$ $10^{6}$
Knowledge Guide Top 1(%) 95.8$\pm$1.0 80.4$\pm$3.9 46.3$\pm$1.2 18.0$\pm$0.8 5.8$\pm$1.7
Top 5(%) 99.8$\pm$0.2 96.8$\pm$0.5 77.8$\pm$1.1 41.6$\pm$1.6 16.6$\pm$2.3
Top 10(%) 100.0$\pm$0.0 98.8$\pm$0.3 87.7$\pm$1.0 53.9$\pm$1.8 25.2$\pm$3.2
Top 25(%) 100.0$\pm$0.0 99.6$\pm$0.2 94.8$\pm$0.6 71.6$\pm$0.6 37.8$\pm$4.1
SP-ID Top 1(%) 95.3$\pm$0.8 78.6$\pm$2.7 35.8$\pm$3.8 12.9$\pm$1.6 3.4$\pm$0.9
Top 5(%) 95.4$\pm$0.1 77.3$\pm$0.7 44.7$\pm$2.3 16.2$\pm$2.4 4.4$\pm$1.5
Top 10(%) 100.0$\pm$0.0 97.3$\pm$0.7 77.5$\pm$2.3 40.2$\pm$2.4 12.3$\pm$1.5
Top 25(%) 100.0$\pm$0.0 99.1$\pm$0.2 85.9$\pm$1.0 53.1$\pm$3.0 18.5$\pm$1.8
WP-ID(th=1) Top 1(%) 92.9$\pm$0.6 71.7$\pm$1.0 32.7$\pm$1.3 10.7$\pm$0.5 3.6$\pm$0.7
Top 5(%) 99.6$\pm$0.1 93.8$\pm$0.8 63.9$\pm$1.5 29.3$\pm$1.5 10.2$\pm$1.2
Top 10(%) 99.9$\pm$0.0 97.1$\pm$0.4 76.8$\pm$0.7 39.3$\pm$0.9 15.7$\pm$1.5
Top 25(%) 100.0$\pm$0.0 99.1$\pm$0.2 88.2$\pm$0.6 55.7$\pm$1.1 26.5$\pm$2.0
No communication Top 1(%) 94.8$\pm$1.2 77.6$\pm$1.4 40.1$\pm$1.2 14.4$\pm$0.9 4.1$\pm$1.1
Top 5(%) 99.8$\pm$0.1 96.2$\pm$0.5 73.6$\pm$2.2 35.8$\pm$1.3 11.4$\pm$1.0
Top 10(%) 99.9$\pm$0.1 98.6$\pm$0.3 84.1$\pm$1.4 47.3$\pm$1.8 17.3$\pm$1.8
Top 25(%) 100.0$\pm$0.0 99.7$\pm$0.2 92.9$\pm$0.9 65.1$\pm$2.2 27.9$\pm$2.5
Appendix D Additional Discussion about Isomers
D.1 Isomer Category
Isomers typically fall into two main categories: constitutional (structural) isomers, which share the same chemical formula but display distinct atom connectivity, and stereoisomers (spatial isomers), which share the same topology graph but diverge in their three-dimensional arrangement (see Figure D.1). Constitutional isomers are NMR-variant, meaning that different isomers produce distinct NMR spectrum. In the sub-categories of stereoisomers, enantiomers are NMR-invariant, but diastereomers and cis-trans isomers are NMR-variant.
Figure D.1: NMR Variability in Isomers
D.2 Isomers Group for $\text{C}_{7}\text{H}_{11}\text{NO}_{3}$
Here is an example for isomer groups. In this isomer group of $\text{C}_{7}\text{H}_{11}\text{NO}_{3}$, they all share the same chemical formula in Figure D.2. The first 10 are constitutional (structural) isomers of each other (cycled green), the last 4 are two pairs of diastereomers (cycled brown). Each of these isomers corresponds to a distinct NMR spectrum.
Figure D.2: Isomer demo for $\text{C}_{7}\text{H}_{11}\text{NO}_{3}$
Appendix E Additional Results on Peak Assignment
In complex natural product molecules, it is a common situation that the local contents of some atoms within the same molecule exhibit a high degree of similarity. It gives rise to challenges for the atomic alignment, as some atoms correspond to ppm values in close proximity. However, our K-M3AID model is capable of recognizing each of the atoms with effective learnt embeddings and deciphering the correspondences among the atoms and the peaks at zero-shot. Two complex natural product molecules with multiple rings (4 and 4, respectively) and multiple chiral centers (6 and 8, respectively) are taken to showcase the effectiveness of atomic alignment (see Appendix Figure E.1).
In molecular A in Figure  E.2, atom 13 and atom 14 are tertiary carbons (attaching to 3 carbons and 1 hydrogen) and on the same 5-member ring, corresponding to the ppm of 34.3 and 35.6, respectively. The similar local content of these two atoms fools SP-ID and WP-ID. In addition, WP-ID fails with more atomic alignments. The molecular B is chemical symmetric regarding atom 0. Thus, atom 1 and atom 3 correspond to the same peak on the spectra. The ppm of atom 1 and atom 3 is 114.2, the ppm of atom 2 and atom 4 is 110.0. While there is 4.2 difference, SP-ID and WP-ID fails to pick up right alignment for atom 1 and atom 3. In contrast, K-ID succeed to align the atoms with peaks in both molecules.
Figure E.1: Examples of Zero-shot Atomic Alignment for Complex Natural Products. Yellow cells in the PPM difference represent the ground truth alignment.
Figure E.2: Extra case studies of IE-Meta-MMA.Yellow cells in the PPM differerence represent the ground truth alignment, and red cross represents the wrong alignment.